1. The age of cars in the staff parking lot of a suburban college is uniformly distributed from six months (0.5) to 11.5 years. Find the 85th percentile

a. 9.85

b.9.15

c. 9

d. 8.5

2. Suppose that the value of stock varies each day from $14 to $25 with a uniform distribution.

Find the 90th percentile

a. 22.1

b. 20

c. 9.9

d. 23.9

3. Find the probability of the shaded region. Round answer to three decimals.

11 10 X 0 1 2 3 4 5 6 7 8 9 10
Answer #1

1)

P(X≥x) =    0.85
(b-x)/(b-a) =    0.15
(11.5- x )/(11.5-0.5)=   0.15
x=   9.85

2)

P(X≥x) =    0.9
(b-x)/(b-a) =    0.1
(25- x )/(25-14)=   0.1
x=   23.9

3)

P (    2.5   ≤ X ≤    5.5   ) =(x2-x1)/(b-a) =    (5.5-2.5)/(10-0)=   0.3

Please let me know in case of any doubt.

Thanks in advance!


Please upvote!